r2 - 09 Feb 2010 - 13:46:46 - KaushikDeYou are here: TWiki >  Admins Web > MinutesDataManageFeb9

MinutesDataManageFeb9

Introduction

Minutes of the US ATLAS Data Management meeting, Feb. 9, 2010
  • Previous meetings and background : IntegrationProgram
  • Coordinates: Tuesdays, Noon Central
    • (605) 715-4900, Access code: 735188; Dial *6 to mute/un-mute.

Attending

  • Meeting attendees:
  • Apologies: Shawn, Bob, Wei, Charles, Saul, John, Wensheng, Hiro, Patrick, Armen
  • Guests:

Topics for this week

  • Space Usage - Armen
    • Allocate more space to MCDISK? Add another 100 TB.
    • Follow up from last week
      • srm monitoring showed 100 TB drop last week - it was due to dcache pool configurations at BNL. Solved.
      • Hiro will provide breakdown by project name - done. Need to study data - for next next week.
      • Check if ESD's generated in the US are being kept on tape - Kaushik.
    • Action item for next week - can we reduce space usage at MWT2? After that NET2? WT2? AGLT2? SWT2?
  • PRODDISK cleaning and consistency - Charles, Hiro
    • DQ2 api change - cleanse_PRODDISK.py was triggering central deletion.
    • Charles released another new version - v*.14.
    • Hiro/Charles will check everything is OK with this version.
    • Please do not run old version - wait till new version is verified.
  • Tier 2 USERDISK cleanup - Hiro, Armen
    • Similar issue as cleanse - cleanup was triggering central deletion.
    • Fixed now - done.
    • Central operations has cleaned up 50k requests.
    • Email went out today to users, for all sites.
  • HOTDISK survey of US ANALY sites - Armen
    • OU is online and missing HOTDISK - make them offline.
    • Duke is offline and missing - send them email.
    • Illinois is offline and missing - Hiro will add.
  • Checksum issue at SLAC - Wei
    • Switched to new checksum program - created a bug. Fixed now.
    • How to clean up mess? Not necessary - all jobs failed. already in Panda.
    • Only affected pilot - not DQ2.
  • Local site mover issue at NET2
    • Found problem with files that have _DQ2* exetension - Athena cannot use them.
    • Paul will update lsm in pilot code.
    • All files will be cleaned up using script from Charles.
    • Charles - periodically, good idea to run Charles script to clean up this debris, and discover what new debris could be collecting (good indicator of file system problems).
  • Lost files on MCTAPE - Pedro
    • Wensheng will follow up and clean Panda and LFC/DQ2.
  • Hot issues
    • SLAC set up SCRATCHDISK and LOCALGROUPDISK - need to be added to SS.
    • Next week - who should be allowed on LOCALGROUPDISK?
    • Wensheng - 3.5 TB used on LOCALGROUPDISK, out of which 2.1 TB are actually official production files. Metadata says custodial (random check of one dataset)! Will check all dataset locations and report next week.
  • AOB


-- KaushikDe - 09 Feb 2010

About This Site

Please note that this site is a content mirror of the BNL US ATLAS TWiki. To edit the content of this page, click the Edit this page button at the top of the page and log in with your US ATLAS computing account name and password.


Attachments

 
Powered by TWiki
This site is powered by the TWiki collaboration platformCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback